OpenStudy (anonymous):

Write in scientific notation. simplify. 4( 2 * 10^-3)

OpenStudy (australopithecus):

8*10^-3

OpenStudy (australopithecus):

\[(\frac{2}{10^3})4 = \frac{(2*4)}{10^3} = \frac{8}{10^3} = 8*10^{-3}\]

OpenStudy (anonymous):

Scientific notation looks like this: \[\large \textrm{a} \cdot 10^\textrm{b}\]Where \(a\) is any number, like \(-13.4, \pi, \sqrt{5.4}\). \(b\) is an integer, like \(-2, -1, 0, 1, 2, 2432, 9325801, \textrm{etc}\). So, we have the \[\large 10^{-3}\] already. What is there left for us to do, then?

OpenStudy (anonymous):

thanksss :)

OpenStudy (anonymous):

Sorry, I just realized I was sort of wrong. \(a\) must be a number like \(0.1, \ 1.234, \ -9.234\) but it CAN NOT be \(12.3, \ \pi, \ \sqrt{5.4}\) Excuse me for the fail. :)

OpenStudy (anonymous):

Basically, \(a\) must begin with only one single digit followed by a decimal point followed by any other numbers.

OpenStudy (anonymous):

Is all gewd! appreciate the explaining :))